# UStyle: Waterbody Style Transfer of Underwater Scenes by Depth-Guided Feature Synthesis

### v2

@@ Abstract
The concept of waterbody style transfer remains largely unexplored in the underwater imaging and vision literature. Traditional image style transfer (STx) methods primarily focus on artistic and photorealistic blending, often failing to preserve object and scene geometry in images captured in high-scattering mediums such as underwater. The wavelength-dependent nonlinear attenuation and depth-dependent backscattering artifacts further complicate learning underwater image STx from unpaired data. This paper introduces UStyle, the first data-driven learning framework for transferring waterbody styles across underwater images without requiring prior reference images or scene information. We propose a novel depth-aware whitening and coloring transform (DA-WCT) mechanism that integrates physics-based waterbody synthesis to ensure perceptually consistent stylization while preserving scene structure. To enhance style transfer quality, we incorporate carefully designed loss functions that guide UStyle to maintain colorfulness, lightness, structural integrity, and frequency-domain characteristics, as well as high-level content in VGG and CLIP (contrastive language-image pretraining) feature spaces. By addressing domain-specific challenges, UStyle provides a robust framework for no-reference underwater image STx, surpassing state-of-the-art (SOTA) methods that rely solely on end-to-end reconstruction loss. Furthermore, we introduce the UF7D dataset, a curated collection of high-resolution underwater images spanning seven distinct waterbody styles, establishing a benchmark to support future research in underwater image STx. The UStyle inference pipeline and UF7D dataset is released at: https://github.com/uf-robopi/UStyle.
Index Terms:
Style Transfer; Waterbody Fusion; Underwater Imaging; Underwater Vision.
@@ I Introduction
Image style transfer (STx) in underwater imagery has significant potential in data augmentation, robotic vision, photometry, and imaging technologies [1]. Traditional neural style transfer (NSTx) approaches [2, 3] primarily focus on terrestrial imagery, aiming to transfer artistic or photometric features across images. However, the concept of “waterbody style” transfer presents a distinct challenge [1], where the goal is to generate an underwater image in the style of a different aquatic environment while preserving the structural integrity of the scene. Although unpaired images from various waterbodies can be collected, conventional NSTx techniques—optimized for perceptual photorealism [4] or artistic blending [5, 6, 7]—struggle to accommodate the unique distortions introduced by underwater light absorption and scattering [8]. As a result, these methods often produce unnatural style adaptation and significant structural loss.
Currently, no learning-based frameworks exist for waterbody STx in underwater imagery. Our previous work, AquaFuse [1], introduced a physics-based approach for waterbody fusion, facilitating data augmentation across different aquatic environments. However, this method relies on a closed-form formulation that requires a carefully selected, noise-free reference image and prior scene information, including depth, incident angle, and attenuation parameters. Consequently, leveraging large-scale vision-based learning models for robust underwater STx without explicit reference images or prior information remains a challenge. Recent advancements in NST [4, 7] and vision-language models [9, 10] present an opportunity to address this gap. However, developing a learning-driven approach capable of accurately adapting underwater distortions while preserving structural details remains an unexplored area of research.
In this paper, we propose UStyle, the first data-driven learning pipeline for waterbody STx in underwater imagery. We present the model architecture, choice of loss functions, and the training strategy required to address the limitations of existing NSTx pipelines for underwater STx. Specifically, UStyle integrates a residual encoder-decoder to extract local structural details at different spatial resolutions while leveraging a progressive multi-stage training strategy to refine waterbody feature learning by domain-specific loss functions. Moreover, to guide the stylization process, we use a physics-based method [1] to synthesize the waterbody from the style image by following the state-of-the-art (SOTA) underwater image formation model [8]. This synthesized waterbody style is then fused by a novel depth-guided whitening and coloring transform (DA-WCT) mechanism, which makes sure that no object/scene information is transferred to the content image. A few sample examples are illustrated in Fig. 1.
@@
Figure 1: Visualizations of the underwater image STx capabilities of UStyle are shown. The content image (top left) is transformed into different waterbody styles. For any given style image, note that its waterbody characteristics are transferred to the content scene, while its object geometry and structural details of the content image remain preserved. Such no-reference waterbody fusion is useful as a geometry-preserving data augmentation tool in underwater imaging, photometry, and robotics applications.
@@
For comprehensive evaluation, we compile an underwater dataset named UF7D – featuring $4,050$ high-resolution images with paired depth annotations across seven different waterbody styles: Blue (B), Green (G), Deep Blue (DB), Deep Green (DG), Clear (C), Blue-Green (B-G), and Green-Blue (G-B). It enables data-driven learning for underwater STx, allowing the model to generalize across diverse aquatic environments with varying optical properties. Most images are sourced from existing SOTA databases [11, 12, 13, 14, 15] for imaging, object detection, image enhancement, and robotics applications; thus, it serves as a useful benchmark for future research. To this end, we conducted a thorough benchmark evaluation of underwater STx, by comparing UStyle with several SOTA methods named StyTR-2 [7], MCCNet [16], ASTMAN [17], IEContraAST [5], PhotoWCT2 [4], and AquaFuse [1].
@@
  1. $1$)
    Underwater STx learning pipeline and dataset: UStyle is the first data-driven STx framework for underwater scenes, eliminating the need for explicit scene priors required by physics-based STx models [1]. Instead of relying on carefully selected reference images and waterbody parameters, UStyle employs a learning-based STx approach that robustly adapts to diverse aquatic environments with varying waterbodies. We also introduce the UF7D dataset, a comprehensive collection of $4,050$ high-resolution underwater images of waterbody styles, which supports supervised training and benchmark evaluation across varying underwater conditions.
  2. $2$)
    Novel contributions in UStyle: We design and develop UStyle to (learn to) blend image features while being aware of the object geometry in underwater scenes for waterbody STx. Our proposed architecture (a) integrates a ResNet-based [18] encoder-decoder with hierarchical skip connections [19] for effective multi-scale feature fusion; (b) employs a progressive multi-stage training strategy [20] for stabilized style-content blending; and (c) domain-specific loss functions learn image statistics (color, lightness) [21], structural details (SSIM [22], FFT [23]) and high-level image contents (VGG [24], CLIP [25]). This guided training strategy is leveraged by our proposed DA-WCT mechanism for depth-aware waterbody stylization, ensuring perceptual fidelity and geometric consistency.
  3. $3$)
    SOTA performance and benchmark evaluation: We conduct comprehensive experimental evaluations validating that UStyle consistently outperforms leading STx methods—such as StyTR-2 [7], MCCNet [16], ASTMAN [17], IEContraAST [5], and PhotoWCT2 [4]—across key metrics such as PSNR, SSIM, RMSE, and LPIPS. Ablation studies also reveal that our proposed DA-WCT mechanism and domain-specific loss functions can deal with the underwater image distortions to generate perceptually coherent stylization. We also compare the performance with physics-based approaches—such as AquaFuse [1]—to validate that UStyle-generated images preserve the structural contents of the scene while fusing waterbody styles. These results establish UStyle as a robust and effective solution for no-reference underwater image STx applications.
@@ II Background and Related Work
Image style transfer (STx) techniques blend the content of one image with the style of another [26, 27, 28], unlocking a wide array of applications across computer vision, virtual reality, robotics, and digital art creation. We can group current advancements in STx into critical domains such as artistic, photorealistic, non-photorealistic, semantic, and multimodal techniques. Each category uses its methods to improve the adaptability and realism of STx, such as optimization algorithms, feed-forward architectures, attention mechanisms, and transformers (see Fig. 2). These approaches collectively push the boundaries of what is achievable in image synthesis, making it possible to preserve structure, adapt to high-resolution outputs, and achieve detailed, context-aware stylization.
@@
  Optimization models [26, 29, 27, 30, 31]
Diffusion-based models [32, 33, 34, 35]
Feed-Forward models [36, 37, 38, 4, 39]
Attention models [40, 41, 42, 43]
Transformers [44, 7, 45, 46, 47, 48, 49]
CNNs and GANs [50, 51, 52, 53, 54]
Physics-based models [55, 56, 57, 58, 59]
Cartoonization models [60, 61, 62, 63, 64]
Texture/pattern analysis [65, 3, 66, 67]
Segmentation models [68, 69, 70, 71, 72, 73]
Region-based models [74, 75, 76, 77]
Multi-style fusion [70, 78, 79, 80]
Few-shot & Zero-shot [81, 82, 83]
Vision-language models [9, 10, 84]
Figure 2: A snapshot of the SOTA literature on image-based style transfer (STx) approaches.
II-A Artistic and Non-Photorealistic Style Transfer
Artistic STx aims to apply one image’s artistic characteristics, such as brushstrokes, textures, or color palettes, to another image’s content. Following the seminal work by Gatys et al. [85], many deep visual learning-based networks have been proposed [36, 37, 38]. In particular, Johnson et al. [86] and Ulyanov et al. [87] develop feed-forward networks that apply predefined styles in real-time, making the process much faster. One limitation of their approach is that their network needs to be trained for a specific style. Huang and Belongie [36] address this issue with adaptive instance normalization (AdaIN). AdaIN adjusts the mean and variance of content features to match style features, enabling the transfer of arbitrary styles [88, 89, 90]. Li et al. [37] introduce the whitening and coloring transform (WCT), which applies a whitening operation to remove content, followed by a coloring operation to inject style statistics.
Non-photorealistic rendering (NPR) methods are a superset of artistic STx, focusing on cartooning, sketches, and more abstract art. Early work in this domain is grounded in traditional computer graphics with handcrafted filters and procedural algorithms, whereas recent work uses deep learning techniques to achieve more sophisticated stylizations. Zhao et al. [60] introduces a GAN-based approach for high-quality image cartoonization, capturing stylistic nuances with remarkable fidelity. Ahn et al. [61] develop an interactive framework that allows users to influence the stylization in real-time, enhancing creativity and personalization. Chen et al. [62] proposes DecoGAN, enabling controllable decoration and stylization of images with user inputs. Additionally, techniques such as language-guided STx [65], fast neural doodling [3], and style consistency models [66] expand the possibilities of NPR, making sophisticated stylization accessible to a broader audience. Kim et al. [67] focuses on texture and pattern analysis to preserve fine-grained details during STx, resulting in more cohesive and visually appealing images. These advancements collectively push the boundaries of NPR, enabling the creation of diverse and artistically rich images through deep learning techniques.
Various attention mechanisms have enhanced artistic STx in recent years by capturing complex relationships between content and style. For instance, self-attention GANs [40] enables different parts of an image to interact more effectively to enhance spatial awareness in the artistic rendering process. Building on this, style-attention networks (SANet) [41] achieve more precise style adjustments in different image areas for refined artistic control. These advancements help balance global and local features, improving the quality of fine-grained stylization and overall composition.
II-B Photorealistic Style Transfer
Photorealistic STx applies stylistic changes while keeping the realism and structure of the original image intact. This is essential in understanding visual scenes and robotics applications. Early methods like deep photo style transfer (DPST) [29] use optimization techniques to achieve photorealism, although they struggle with high-resolution input images. To address this, Li et al. introduces PhotoWCT [38], which uses cascaded autoencoders to transform content features while keeping structural details intact. PhotoWCT2 [4] improves this by reducing parameters with blockwise training for faster stylization of high-resolution images. Wavelet-based WCT2 [39] uses wavelet pooling and un-pooling layers to preserve fine details during the STx process better.
Beyond feed-forward models, vision transformers excel at modeling long-range dependencies and spatial coherence, making them ideal for tasks where preserving local details and global structure is essential. StyTr2 by Deng et al. [7] is a prominent work that employs dual transformer encoders: one each for content and style; it introduces the idea of content-aware positional encoding (CAPE) to ensure scale invariance and spatial preservation. Contemporary works also explore unsupervised STx learning, with models like TransGAN [50], integrating transformers with GANs. More advanced attention mechanisms such as transformer-in-transformer (TNT) [48] further enhance STx by capturing complex dependencies through a nested architecture. Recent developments in diffusion-based models [32, 35] demonstrate further improvements by disentangling style and content for controlled (localized) adjustments.
On the other hand, physics-based methods use precise rendering to improve photorealism. For instance, Monte Carlo ray tracing simulates realistic lighting and shading to capture natural light interactions. Bidirectional reflectance distribution function (BRDF) modeling [57] preserves material properties like glossiness and roughness for more realistic appearance. These models consider how light reflects off surfaces to capture various light-material interactions. Another notable approach is deep image prior (DIP) [56], which uses a neural network to learn photorealism directly from the input image. It leverages natural image priors (  e.g., smooth regions, sharp edges) to handle tasks such as denoising, restoration, and style transfer. Moreover, retinex theory [91] powers many algorithms to improve photorealism by exploiting the geometry of how the human eye separates light from surface reflectance. Techniques like multi-scale retinex [92] restore true colors by ensuring natural color balance and contrast. Inverse tone mapping techniques [55] focus on rendering high-dynamic-range images, whereas optical flow-based STx [58] maintain spatial coherence to ensure smooth transitions in motion, shading, and texture gradients.
II-C Semantic and Multimodal Style Transfer
Semantic STx integrates a high-level understanding of image content to apply styles context-awarely, enhancing the coherence between the transferred style and the underlying content. For instance, Wang et al. [72] employs the SOTA semantic segmentation model DeepLab [71] to preserve object boundaries during STx. Lin et al. [68] further refine this approach by leveraging segmentation to maintain structural details, ensuring context-aware consistent STx. On the other hand, region-based models focus on selectively transferring styles to specific areas within an image. Niu et al. [74] propose a region-adaptive STx method that applies styles to distinct regions based on their semantic relevance. Liao et al. [77] develop a semantic-guided attention mechanism to enhance stylization in targeted regions, improving aesthetic quality and content preservation. These approaches allow for more precise control over the stylization process by fusing global semantic context and local features for improved STx.
Multimodal STx extends these concepts by incorporating additional modalities, such as depth information or textual descriptions, to guide the stylization. Ye et al. [70] explores multimodal learning techniques to fuse visual and textual data, enabling more expressive and versatile STx. Zheng et al. [69] applies retinal image segmentation combined with STx for medical imaging applications, demonstrating the versatility of multimodal approaches. Besides, few-shot and zero-shot learning techniques have been introduced to address the challenge of limited-style exemplars. Fu et al. [81] present StyleAdv, a few-shot adversarial learning framework that can adapt to new styles with minimal data. Huang et al. [82] proposes a zero-shot STx method using residual decay networks, eliminating the need for style images during training. Liu et al. [83] introduce StyleRF, which leverages neural radiance fields for zero-shot STx, enabling the application of novel styles without prior exposure.
Vision-language models (VLMs), particularly Contrastive Language–Image Pre-training (CLIP), enable the integration of textual and visual information for advanced image processing tasks [84]. Their robust image encoders capture comprehensive representations of image features, enhancing the potential for applications such as image STx. Clipstyler [9] framework achieves image STx solely from text descriptions by leveraging a patch-wise CLIP loss and perspective augmentation, thus eliminating the need for reference (style) images. Yang et al. [10] propose a zero-shot contrastive loss for diffusion models [33] that leverage patch-wise contrastive learning to preserve semantic content during text-guided STx without additional fine-tuning or auxiliary networks. These advancements significantly enhance the flexibility and scalability of STx models, making them more practical for real-world applications.

### v3

@@ Abstract
The concept of “waterbody style” transfer remains largely unexplored in the underwater imaging literature. Traditional image style transfer (STx) methods emphasize artistic and photorealistic blending that fails to preserve geometry in high-scattering underwater environments. The wavelength-dependent attenuation and depth-dependent backscattering of underwater optics further complicate STx learning from unpaired data. We introduce UStyle, the first data-driven framework for transferring waterbody styles across underwater images without needing reference images or explicit scene information. We propose a novel depth-aware whitening and coloring transform (DA-WCT) that incorporates physics-based waterbody synthesis for perceptually consistent stylization while preserving scene structure. We also integrate carefully designed loss functions to maintain color, lightness, structural integrity, frequency-domain features, and high-level content in VGG and CLIP spaces. Comprehensive experimental analyses show that UStyle surpasses state-of-the-art methods that rely solely on reconstruction loss. Additionally, we present the UF7D dataset, a curated benchmark of high-resolution underwater images across seven waterbody styles. The UStyle code and UF7D dataset are available at: https://github.com/uf-robopi/UStyle.
@@ I Introduction
Image style transfer (STx) in underwater imagery has significant potential in data augmentation, robotic vision, photometry, and imaging technologies [siddique2024aquafuse]. Traditional neural style transfer (NSTx) approaches [chen2024upst] primarily focus on terrestrial imagery, aiming to transfer artistic or photometric features across images. However, the concept of “waterbody style” transfer presents a distinct challenge [siddique2024aquafuse], where the goal is to generate an underwater image in the style of a different aquatic environment while preserving the structural integrity of the scene. Although unpaired images from various waterbodies can be collected, conventional NSTx techniques—optimized for perceptual photorealism [chiu2022photowct2] or artistic blending [chen2021artistic, deng2021arbitrary, deng2022stytr2]—struggle to accommodate the unique distortions introduced by underwater light absorption and scattering [akkaynak2018revised].
@@
Figure 1: Visualizations of UStyle’s underwater STx capabilities are shown. The content image (top left) is transformed into various waterbody styles to transfer the waterbody characteristics while preserving object geometry and structural details.
Currently, no learning-based frameworks exist for waterbody STx in underwater imagery. Our previous work, AquaFuse [siddique2024aquafuse], introduced a physics-based approach for waterbody fusion, facilitating data augmentation across different aquatic environments. However, this method relies on a closed-form formulation that requires a carefully selected, noise-free reference image and prior scene information, including depth, incident angle, and attenuation parameters. Leveraging large-scale vision-based learning models for robust underwater STx without reference images or prior information remains a challenge. Recent advancements in NST [chiu2022photowct2, deng2022stytr2] and vision-language models [kwon2022clipstyler, yang2023zero] present an opportunity to address this gap.
In this paper, we propose UStyle, the first data-driven learning pipeline for waterbody STx in underwater imagery. We present the model architecture, choice of loss functions, and the training strategy to address the limitations of existing NSTx pipelines. UStyle integrates a residual encoder-decoder to extract local structural details at different spatial resolutions while leveraging progressive multi-stage training to refine waterbody features by domain-specific loss functions. To guide the stylization process, we use a physics-based method to synthesize the waterbody from style image by following the state-of-the-art (SOTA) underwater image formation model [akkaynak2018revised]. This synthesized waterbody style is then fused by a novel depth-guided whitening and coloring transform (DA-WCT) mechanism, which makes sure that no object texture or scene information is transferred to the content image. A few sample examples are illustrated in Fig. 1.
@@
For comprehensive evaluation, we compile an underwater dataset named UF7D – featuring $4,050$ high-resolution images with paired depth annotations across seven different waterbody styles: Blue (B), Green (G), Deep Blue (DB), Deep Green (DG), Clear (C), Blue-Green (B-G), and Green-Blue (G-B). It enables data-driven learning for underwater STx, allowing the model to generalize across diverse aquatic environments with varying optical properties. About 5858% of the images are sourced from existing SOTA databases [li2019underwater, akkaynak2019sea, peng2023u, islam2020fast] for imaging, object detection, image enhancement, and robotics applications; thus, it serves as a useful benchmark for future research. The remaining $42\%$ data are collected from a diverse set of open-access sources, where we solely focused on ensuring diversity of object-scenes geometry and waterbody types. We also release a UF7D-Challenge test set with $70$ samples across categories for more challenging evaluation of underwater STx methods. To this end, we conducted a thorough benchmark evaluation of underwater STx, by comparing UStyle with several learning-based as well as physics-based SOTA methods named StyTR-2 [deng2022stytr2], MCCNet [wang2020arbitrary], ASTMAN [deng2020arbitrary], IEContraAST [chen2021artistic], StyleID [chung2024style], PhotoWCT2 [chiu2022photowct2], and AquaFuse [siddique2024aquafuse].
@@
  1. 11)
    Underwater STx learning pipeline and dataset: UStyle is the first data-driven STx framework for underwater scenes, eliminating the need for explicit scene priors required by physics-based STx models [siddique2024aquafuse]. Instead of relying on carefully selected reference images and waterbody parameters, UStyle employs a learning-based STx approach that robustly adapts to diverse aquatic environments with varying waterbodies. We also introduce the UF7D dataset, a comprehensive collection of $4,050$ high-resolution underwater images of waterbody styles, which supports supervised training and benchmark evaluation across varying underwater conditions.
  2. 22)
    Novel contributions in UStyle: We design and develop UStyle to (learn to) blend image features while being aware of the object geometry in underwater scenes for waterbody STx. Our proposed architecture (a) integrates a ResNet-based [koonce2021resnet] encoder-decoder with hierarchical skip connections [seo2019hierarchical] for effective multi-scale feature fusion; (b) employs a progressive multi-stage training strategy [wang2018progressive] for stabilized style-content blending; and (c) domain-specific loss functions learn image statistics (color, lightness) [dong2022underwater], structural details (SSIM [hore2010image], FFT [brigham1988fast]) and high-level image contents (VGG [simonyan2014very], CLIP [radford2021learning]). This guided training strategy is leveraged by our proposed DA-WCT mechanism for depth-aware waterbody stylization, ensuring perceptual fidelity and geometric consistency.
  3. 33)
    SOTA performance and benchmark evaluation: We conduct comprehensive experimental evaluations validating that UStyle consistently outperforms leading STx methods such as StyTR-2 [deng2022stytr2], MCCNet [wang2020arbitrary], ASTMAN [deng2020arbitrary], IEContraAST [chen2021artistic], PhotoWCT2 [chiu2022photowct2], and StyleID [chung2024style] across key metrics such as PSNR, SSIM, RMSE, and LPIPS. Ablation studies over all loss components also reveal that our proposed DA-WCT mechanism and domain-specific loss functions can deal with the underwater image distortions to generate perceptually coherent stylization. We also compare the performance with physics-based approaches such as AquaFuse [siddique2024aquafuse] – to validate that UStyle-generated images preserve the structural contents of the scene while fusing waterbody styles. These results establish UStyle as a robust and effective solution for no-reference underwater image STx applications.
@@ II Background and Related Work
Image style transfer (STx) methods combine the content of one image with the style of another [gatys2015neural, liu2024swinit], enabling applications in computer vision, AR/VR, robotics, and digital art. Existing STx methods include artistic, photorealistic, semantic, and multimodal techniques that employ optimization, feed-forward networks, attention mechanisms, and transformers; see Fig. 2.
II-A Artistic and Non-Photorealistic STx
Artistic STx blends one image’s artistic characteristics, such as brushstrokes, textures, or color palettes, into another image’s content. Following the seminal work by Gatys et al. [jing2019neural], many deep visual learning-based networks have been proposed [huang2017arbitrary, li2017universal, li2018closed]. In particular, Johnson et al. [johnson2016perceptual] and Ulyanov et al. [ulyanov2016instance] develop feed-forward networks that apply predefined styles in real-time, making the process much faster. One limitation of their approach is that their network needs to be trained for a specific style. Huang and Belongie [huang2017arbitrary] address this issue with adaptive instance normalization (AdaIN), enabling the transfer of arbitrary styles [chen2022quality, wang2022clast, chen2022quality]. Li et al. [li2017universal] introduce the whitening and coloring transform (WCT), which applies a whitening operation to remove content, followed by a coloring operation to inject style statistics. A superset of artistic STx is Non-photorealistic rendering (NPR) that focuses on cartooning, sketches, and more abstract art [ahn2023interactive]. Ahn et al. [ahn2023interactive] develop an interactive framework that allows users to influence the stylization in real-time, enhancing creativity and personalization [chen2024decogan]. Additionally, techniques such as language-guided STx [fu2022language], and style consistency models [zhang2023style] expand the possibilities of NPR, making sophisticated stylization accessible to a broader audience [kim2019transport].
@@ II-A Artistic and Non-Photorealistic STx
  Optimization models [gatys2015neural, luan2017deep, wang2019research, kim2020deformable, lyu2023infostyler]
Diffusion-based models [zhang2023inversion, wang2023stylediffusion, chung2024style, li2024diffstyler]
Feed-Forward models [huang2017arbitrary, li2017universal, li2018closed, chiu2022photowct2, yoo2019photorealistic]
Attention models [zhang2019self, park2019arbitrary, liu2021adaattn, wang2022esa]
Transformers [vaswani2017attention, deng2022stytr2, liang2021swinir, wang2022uformer, chen2022domain, han2021transformer, fengxue2023image]
CNNs and GANs [jiang2021transgan, fu2019dual, lin2020gan, qian2022circular, xu2021learning]
Physics-based models [banterle2009inverse, ulyanov2018deep, lafortune1997non, horn1981determining, he2010single]
Cartoonization models [zhao2024gan, ahn2023interactive, chen2024decogan, li2019everyone, wu2021deep]
Texture/pattern analysis [fu2022language, zheng2024fast, zhang2023style, kim2019transport]
Segmentation models [lin2021image, zheng2023retinal, ye2023multi, chen2017deeplab, wang2020image, madokoro2022semantic]
Region-based models [niu2021style, yamaguchi2015region, krutarth2021region, liao2022semantic]
Multi-style fusion [ye2023multi, huang2019style, zhang2018multi, yu2024multi]
Few-shot & Zero-shot [fu2023styleadv, huang2020rd, liu2023stylerf]
Vision-language models [kwon2022clipstyler, yang2023zero, cai2023image]
Figure 2: A categorization of the SOTA literature on image-based STx and some prominent methods.
II-B Photorealistic STx
Photorealistic STx applies stylistic changes while preserving an image’s realism and structure, which is essential for scene understanding and robotics. Early methods such as DPST [luan2017deep] use optimization but struggle with high resolutions. To address this, PhotoWCT [li2018closed] employs cascaded autoencoders to maintain structure, while PhotoWCT2 [chiu2022photowct2] and wavelet-based WCT2 [yoo2019photorealistic] improve efficiency and detail preservation. Vision transformers also enhance photorealistic STx; for example, StyTr2 [deng2022stytr2] uses dual transformer encoders with content-aware positional encoding (CAPE) for spatial coherence. Unsupervised methods like TransGAN [jiang2021transgan] and advanced architectures such as TNT [han2021transformer], along with diffusion-based models [zhang2023inversion, li2024diffstyler], further disentangle style and content for controlled adjustments. Additionally, methods like DIP [ulyanov2018deep] leverage natural image priors to restore true colors and preserve spatial coherence.
II-C Semantic and Multimodal STx
Semantic STx integrates a high-level understanding of image content to apply styles context-awarely, enhancing the coherence between the transferred style and the underlying content. For instance, Wang et al. [wang2020image] use DeepLab [chen2017deeplab] to preserve object boundaries, while Lin et al. [lin2021image] refine this approach with segmentation to maintain structure. Region-based method such as Liao et al. [liao2022semantic] enables selective stylization by fusing global semantic context with local features. Few-shot and zero-shot approaches, including StyleAdv [fu2023styleadv], Huang et al. [huang2020rd], and StyleRF [liu2023stylerf], further address the challenge of limited style exemplars. Vision-language models, particularly CLIP [cai2023image], integrate textual and visual information to enhance STx. Clipstyler [kwon2022clipstyler] leverages a patch-wise CLIP loss for text-guided stylization, and Yang et al. [yang2023zero] propose a zero-shot contrastive loss for diffusion models [wang2023stylediffusion] that preserves semantic content without fine-tuning. These advancements significantly improve the flexibility and scalability of STx for real-world applications.